Average Rate of Return—Cost Savings

Master Fab Inc. is considering an investment in equipment that will replace direct labor. The equipment has a cost of $131,000 with a $11,000 residual value and a 10-year life. The equipment will replace one employee who has an average wage of $31,140 per year. In addition, the equipment will have operating and energy costs of $6,360 per year.

Determine the average rate of return on the equipment, giving effect to straight-line depreciation on the investment. If required, round to the nearest whole percent.
%
Answer #1

Cost of equipment = $131,000

Residual value = $11,000

Useful life = 10 years

Annual depreciation = (Cost of equipment - Residual value)/Useful life

= (131,000 - 11,000)/10

= $12,000

Annual saving in labor cost = $31,140

Increase in operating cost = $6,360

Annual saving in labor cost 31,140
Increase in operating cost -6,360
Annual depreciation -12,000
Net income $12,780

Average investment = (Cost of equipment + Salvage value)/2

= (131,000 + 11,000)/2

= $71,000

Average rate of return = Net income/Average investment

= 12,780/71,000

= 18%

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.